OpenStudy (anonymous):

]The formula for degrees Celsius is: c = 5/9 (F − 32), where F stands for degrees Fahrenheit. Which of the following equations shows the formula correctly solved for F?

OpenStudy (anonymous):

f = 5/9c+32 f=9/5c+32 f=5/9c-32 f=9/5c-32

OpenStudy (anonymous):

At first, we can see that 5/9 would have been inverted, so it needs to be 9/5, that eliminates equations 1 and 3. Then we have to decide between -32 or 32, but since it is -32 in the original formula, then when it goes to the other side of the equation it becomes +32, so it would be equation 2